# Experimental evidence of quantum interferences in CO-H2 rotational energy transfer at room temperature

### v1

@@ VI Summary and conclusion
Experimental evidence is reported for the first time for odd and even propensity rules for the CO-H2 system. This represents a valuable and sensitive tool for benchmarking the anisotropy of the PES. This study shows that the CO-H2 system behaves in a significantly different way from the CO-He system from the perspective of both the magnitude and general trends of rate coefficients, including the propensity rules. These findings were made possible by the use of the time-resolved IR-VUV double resonance technique which is able to reveal quantum effects even in thermally averaged experiments thanks to high quantum-state selectivity.

### v2

@@ VI Summary and conclusion
Experimental evidence is reported for the first time for odd and even propensity rules for the CO-H2 system. This represents a valuable and sensitive tool for benchmarking the anisotropy of the PES. This study also shows that the CO-H2 system behaves in a significantly different way from the CO-He system from the perspective of both the magnitude and general trends of rate coefficients, including the propensity rules. These findings were made possible by the use of the time-resolved IR-VUV double resonance technique which is able to reveal quantum effects even in thermally averaged experiments thanks to high quantum-state selectivity. Finally, in future work, a more detailed theoretical investigation would be valuable (e.g. a partial-wave analysis), for a more comprehensive interpretation of the interference phenomenon in molecular scattering.